How many liters of 20% alcohol solution should be added to 40 liters of a 50% alcohol solution to make a 30% solution?
Answer by monika_p(71)

x- liters of 20% alc. solution
(x+40) - total liters of 30% solution
x*20%+40*50%=(x+40)*30%
0.2x+40*0.5=(x+40)*0.3
0.2x+20=0.3x+12
0.1x=8
x=80
You have to add 80 liters of 20% alc. solution